A right triangle has: hypotenuse of 16 cm, side of 6√3 cm. Find the other side.
Answer by mananth(16946)

hyp^2-side^2=side^2
16^2-6sqrt3 ^2 = side ^2
256-36*3= side ^2
256-108= side ^2
48=side^2
sqrt 2*2*2*2*3=side^2
4sqrt3 = side
Ananth